# The unified cross-disciplinary model of the operation of neurons

### v2

@@ 2.8 The dynamic layers
This layer’s importance and operation are in line with the Earth’s atmosphere. Its features drastically deviate from the features of the bulks on its two sides. It is separated by a sharp contour on one side and an ill-defined border on the other; its volume is far from being homogeneous. Gravity keeps it in place, and it is at rest. However, sometimes, for some periods, also other (thermodynamical and electrical) forces are evoked inside it and lead to transient changes, moving huge masses with high speeds inside it. Their thickness and mass are negligible compared to those of the bulks on their two sides, and we can describe the bulks without considering their density, mass, size, etc. Still, this thin layer is responsible for the weather. It can temporarily absorb products of slow processes (water evaporation) and deliver masses of high density (much above its density, such as water, sand, etc.) to continental distances, creating the illusion that it stores that matter. Minor changes (natural ones, such as a slight difference in air temperature, and artificial ones, such as injecting condensation nuclei in clouds) can result in enormous changes. We can even imagine volcanic eruptions as semipermeable gates for material with apparently random operation and distribution of the injected material.
@@
Similar is the case with the neuronal membranes and the neuronal operation. Now, we are at the point where their decades-old static description is insufficient. We need to derive the corresponding laws of motion to describe the neuron’s dynamic behavior. We need a meticulous and unusual analysis to derive them. In a neuron, in the abstraction science uses, we put together only an ionic solution, a semipermeable membrane, and currents that enter and leave it. All these belong to non-living matter. As experienced, at some combination of their parameters and gradients, qualitatively different phenomena happen, which, in the abstraction biology uses, are called signs of life; our system starts to belong to living matter. Given that the approximations, the derived abstractions, and the mathematical formalisms describing them are different for the two cases, it looks like we have two different, only loosely bound worlds. We realize we have arrived at the boundary of non-living and living matters, and we must go back to the first principles of science to clarify where their boundary is. However, by using our approach, we may defy that "the emergence of life cannot be predicted by the laws of physics" [48]. Our artificial duck looks, quacks, and swims like a duck.
@@ 2.9 Ionic currents in a cell
Notice that if any of the factors is zero, the macroscopic current $I$ is zero. Microscopic carriers must be present in the volume, and have charge, the cross section must not be zero, and the charge carriers must move with a potential-assisted speed, which needs an electric field generated by the system, see Eq.(7). When there is no constraint force, the thermodynamic and electrical forces will result in an effective force moving the ions. That is, the ions can move without an external electric field, as experienced in biological systems. One of the fundamental mistakes by HH was neglecting the Coulomb force for ions’ electric interaction, and that the electrical potential is created by diffusion processes instead of ideal electric batteries.
In biological systems, where ions represent the "current", it is either a native current (without an external potential) or an artificially injected current or potential generating a current. This way, the current is consistently accompanied by a change in concentration gradient, as the moving ions represent both mass transfer and charge transfer simultaneously. The potential and current are connected through the features of the medium (material) that hosts our measurement.
@@
(drag force) acting on it. A (microscopic) electric force field $\frac{dV}{dz}$ inside the wire would accelerate the charge carriers continuously
@@
with a constant speed $v$. It is not the drift speed because of the electric repulsion, it is a potential-assisted or potential-accelerated speed that can be by orders of magnitude higher. The medium in which the charge moves shows a (macroscopic, speed-dependent) counterforce $F_{d}$, which in steady state equals $F_{electric}$, that is :
@@
The amount of current in a wire is not only influenced by the electric force field (specific resistance) but also by the number of charge carriers $n$. While the latter is commonly considered constant and part of the former, this is not necessarily the case for biological systems with electrically active structures inside. The medium’s internal structure introduces significant modifications (according to E. Schrödinger: "because the construction is different from anything we have yet tested in the physical laboratory" [22]). Applying an additional electric field to a wire can generate varying amounts of current as the number of charge carriers changes due to biological effects. For axons, we use a single-degree-of-freedom system, a viscous damping model, so the ions will move with a field-dependent constant velocity in the electric space; so it takes time while they appear on the membrane. The activity of potential-controlled ion channels in its wall may change $n$ in various ways; furthermore, that change can result in ’delayed’ currents during measurement, for example, in clamping, as physiological measurements witness.
@@
The higher the resulting space derivative (gradient) and the fewer ions that can share the task of providing a current, the higher the speed. We hypothesized (it needs a detailed simulation) that in the case of this charged fluid, the electric repulsion plays the role of ’viscosity’ [8]. The higher the charge density, the stronger the force equalizing the potential; so $\eta$ is the lower, the higher the charge density (proportional to $C_{k}$). For the sake of simplicity, we assume that the speed is proportional to the space gradient of the local concentration and potential. Recall that our equations refer to local gradients only. The electric gradient can propagate only with the speed of the concentration gradient, given that only the chemically moved ions can mediate the electric field [8]. The lower interaction speed limits the other interaction speed if the interactions generate each other.
@@
It is important to remember for alternating current experiments: the different ions will move with different speeds under the effect of the same $E(z)$, which is the function of the local concentrations. The conduction speed sensitively depends on the temperature, and through it, the shape parameters [50] of the AP and especially the length of the "relative refractory" period [51].
@@ 2.10 Imitating slow currents
At the dawn of finding methods for describing neuronal operation, HH published high-precision measurements [49] enabling detailed testing of theories explaining the observed physiological behavior. Their good physical model that “movement of any charged particle in the membrane should contribute to the total current” lacked considering the mutual repulsion and finite speed of those particles (they are about 50,000 times heavier and a million times slower than the electrons and they move in a closed space segment). Furthermore, they have started from the commonly used wrong assumption that conductance is a primary electric entity. They could "find equations which describe the conductances with reasonable accuracy and are sufficiently simple for theoretical calculation of the AP and refractory period". The "reasonable accuracy" and the need for a simple calculation method (performed on a mechanical calculator) obscured that the laws for electrons are not the same as the laws derived for ions.
@@
There is no voltage-dependent conductance [52]. Instead, the finite speed of ions and the wrong assumption that conductance is a primary entity misleads physiological research. With wording that "conductance changes", one states that charge carriers appear/disappear/reappear; that is, the charge conservation as well as mass conservation are not fulfilled. Claims such as "the current cannot disappear, it has to go somewhere" [25], page 9, are wrong: there is no conservation law for current, which is the derivative of charge with respect to time. The correct physics background behind of the early phenomenon [31] is that the number of charge carriers, at the cross section where the measurement is carried out, changes as wave of ions ions with variable intensity moves slowly along the axon (they are apparently “created” for the static measurement arrangement). The changing conductance is a fallacy stemming from the idea of "fast current" which is valid for electrons but not for slow ions.
@@
During the regular operation of a neuronal membrane, after opening the ion channels, a vast amount of ions flows into the intracellular space from the extracellular space, imitating the effect of switching a clamping voltage. The essential difference is that the ions arrive through the axon to the joining point in clamping. In contrast, through the membrane’s ion channels, they directly contribute on the membrane’s entire surface. The membrane’s size is finite, so with a finite current speed, it takes time until the charges on the membrane’s surface arrive at the AIS, as we discussed for the axonal current. Fig. 5 shows the measured rising and falling edges of the clamping experiment performed by HH [49]. The continuous lines show the diagram lines fitted to the experimental data (notice again that the fitted polynomial significantly differs from the correct exponential function around the time 0.5 msec); the dashed line shows the expected result of the rising edge, calculated with the time constant of the falling edge. The number of charge carriers follows a saturating curve, that decreases the current in the axon. HH measured that their time constants differ far outside their experimental error, that would not be possible if they were the result of charge-up and discharge of the same oscillator circuit. The measured rising edge corresponds to a charge-up process where the charge-up current is a saturating curve due to the change in the number of charge carriers in the axon; see Eq.(23).
Figure 5: The measured current, instead of conductance, (Fig. 2 from [49]), demonstrating that two different physical effects result the rising and falling edge of the measured current.
@@ 2.11 Effects of membrane’s electricity
2.11.1 Maxwell demon
Nature builds ’Maxwell-demon’-like objects into the separating membrane: gated ion channels, see Fig. (4). The ion channels operate as demons (from the point of view of the segments and the observer). In the classical theory, some power opens them, and they autonomously transfer ions in a potential-accelerated operating mode, and then that power puts the cap back on the top of the channel. The segments are separated by a semipermeable membrane having capped ion channels inserted into its material. As long as the caps of the ion channels are closed and the ion concentration on the two sides of the membrane is the same, we do not see any change in the state of the solutions. Although the channels can stochastically open, close, and reopen, they transmit more or less well-defined charge quanta. Even the channels can recognize the ions’ chemical nature and transmit only a selected ion type. The channels are passive during those processes, although the enormous voltage gradient can rearrange their structure and change their behavior through that. The demons also coordinate their actions using the layer containing charges as a communication medium; their population maintains a well-defined macroscopic current across the membrane.
2.11.2 Passing through ion channels
@@
In Fig. 4, an ion channel is depicted in the middle of the figure with a diameter of about 3nm3\ nm for visibility. Furthermore, we assume that the ion’s size and, correspondingly, the thickness of the atomic layer in the electrolyte on the surface of the membrane is approximately 0.1nm0.1\ nm (although with their electric double layers, they can also reach a size of 1nm1\ nm). For comparison, recall that an Atomic Force Microscope (AFM) tip’s size is typically about 10nm10\ nm, the size of the tip of the clamp pipette is in the range of 1,000nm1,000\ nm and the size of the soma in the range of 10,000nm10,000\ nm.
@@
Figure 4 also hints at how ions can move in the proximity of the membrane. On the left side, the electric field increases toward the membrane, so the ions cannot move in that direction without an initial force being applied. The case is the same on the right side because of the opposite charge. An ion must gain energy to get closer to the membrane, which means, at the same time, higher potential energy. This situation explains why the ions do not diffuse into the bulk region. At their present positions, the counterforces by the membrane balance the attraction from the opposite side. The ion channels can work in continuous mode (the "resting ion channels") and impulse mode (the "transient ion channels" are gated by "caps"): "Two types of ion channels—resting and gated— have distinctive roles in neuronal signaling. Resting channels are primarily important in maintaining the resting membrane potential, the electrical potential across the membrane in the absence of signaling. Some types of resting channels are constitutively open and are not gated by changes in membrane voltage; other types are gated by voltage but can open at the negative resting potential of neurons. Most voltage-gated channels, in contrast, are closed when the membrane is at rest and require membrane depolarization to open.[1], page 126.
@@
Notice that the resulting electric field cannot accelerate an ion except when the ion is at the beginning of the open ion channel across the membrane. Here, the mechanical counterforce is missing, and the field drops to the value generated by the condenser inside the plates. Here, a vast electric field accelerates the ions to a high speed (although the force needed to move against the viscosity decreases the electrostatic acceleration). The empty red circles show an ion that traverses through the channel. The figure shows the electric fields on both sides of the channel at the beginning and end. (On the right side, the electric field is negative, but so is the charge of the ions.)
@@
The electric acceleration sharply decreases after leaving the ion channel (see the right side). As shown by the difference in the vertical positions of the filled red and green balls, the difference in the electric field values is huge across the membrane ($E_{accelerate}$) and considerable inside the dielectric electrolyte ($E_{assist}$) as shown by the difference in the electric field values of the green balls. The field still accelerates, but significantly less intensely than the one in the gap. Suppose we assume that the ion quickly accelerates to its Stokes-Einstein speed. In that case, the travel speeds in those space regions are proportional to the difference of the fields in the shown positions and the reciprocals of their travel times.
@@
The passed ions decrease the concentration and the electric field on the high-concentration side, and, conversely, increase them on the low-concentration side. At the exit, the field continues to accelerate the ion. However, the increasing potential and concentration increasingly decelerate the passing ions, so they quickly brake them to the assisted speed. Then, the effect of the electric field cancels, and the ions will move with their diffusion speed ("the ion stops"). The local potentials, connected to the operation of the ion channels, act as an "electric lock".
@@
We note that moving charge inside the ion channels (nanopores) in membranes is a complex phenomenon and its details are discussed in [19]. They have studied theoretically and experimentally the generation of membrane potential at zero current in nanoporous membranes separating two reservoirs with different electrolyte concentrations. Although not specifically for our environment, they confirmed our statement that the membrane potential is of electrostatic origin and in the order of a couple of $mV$. Their finding that "the corresponding Donnan potentials appear at the pore entrance and exit, leading to a dramatic enhancement of membrane potential in comparison with an uncharged dielectric membrane" inspires beginning further research in uncovering quantitatively the details of the operation of the dynamic layer governing neuronal operation; furthermore, moving toward studying non-equilibrium processes.
@@
We want to describe the change of voltage due to limited resources using an idea similar to the predator-prey model. We have charged sheets (represented by ions in the electrolyte) with potential $U_{H}$ and $U_{L}$ on the two surfaces of the membrane, plus we consider the corresponding neighboring layers on their side toward the "bulk" of the segment. We assume that the layers’ capacity is constant and has the same value, so the change in charge is linearly proportional to the change inf the voltage.
@@
We consider that a charge transfer happens from the layer with potential $U_{H}$ to the layer with potential $U_{L}$  (that is, the same charge is removed and added, respectively) with a high speed (we call it potential-accelerated speed); furthermore, that with a much lower speed (we call it potential-assisted speed) the charge in the neighboring layer increases, and decreases, respectively, and those voltages are ($U_{H-1}$ and $U_{L+1}$).
@@
We assume a constant capacity for the layers and that the amount of transferred charge (or voltage) is proportional to the difference in voltages in the respective layers.
@@
As Fig. 7 depicts, voltages $U_{H}$ and $U_{L}$ quickly approach their balanced values. When they get equal, the driving force cancels. In the meantime, the voltages $U_{H-1}$ and $U_{L+1}$ tend to approach $U_{H}$ and $U_{L}$, respectively. The diffusional and energy-producing processes, depending on the local conditions, change the voltages on the two ends of the channel, creating the illusion that the channel opens and closes (ad-hoc constants are used to illustrate the concept). As [27] analyzed, this on/off behavior of currents can be observed for natural and synthetic lipid membranes, without caps. Despite slightly different experimental conditions, the amplitudes and typical time scales are similar.
2.11.3 Ion selectivity
The mechanism of channel passing can also contribute to explaining ion selectivity. "The normal selectivity cannot be explained by pore size, because $Na^{+}$ is smaller than $K^{+}$ [45]". The fact that ion selectivity changes with the membrane potential is observed (under the name of changed specific permeability), but it remains unexplained. "For example, the negative resting potential of nerve cells is largely determined by a class of $K^{+}$ channels that are 100-fold more permeable to $K^{+}$ than to $Na^{+}$. In contrast, during the action potential a class of $Na^{+}$ channels is activated that are 10- to 20-fold more permeable to $Na^{+}$ than to $K^{+}$. Thus a key to the great versatility of neuronal signaling is the regulated activation of different classes of ion channels, each of which is selective for specific ions." [1], page 101. However, the question of why the permeability changes and the channels are activated remains open. It is remarkable that an ion-specific channel is permeable also for other ions, only different transmission intensity is observed.
The commonly used picture of the operation of selectivity filters is surely wrong. The assumed mechanical operation of the pores is too slow: the assumed structural change needs $10^{-8}\ s$ and the ions passage time is about $10^{-11}\ s$ (furthermore, it must be repeated about $10^{3}$ times per passage). If a wrong ion is caught, it must be transported back to its departure side, through the right ions (against their repulsion), and the right ions must retry. Neither moving forward nor backward is an appropriate driving force present in that picture. Furthermore, the transport capacity is not sufficient for the operation.
In our model, as Eq.(6) states, the electrolyte separated by a permeable membrane is balanced when the electrical and thermodynamic forces are equal. The $E_{thermal}^{C_{k}}(d)$ "electric field" depends on the chemical quality of the ion; the electrical field does not. That means the resulting driving forces acting on the different ions are different. While a resulting force is exerted on one ion and moves it to the other segment, the other remains resting, or even moves in the opposite direction. Furthermore, the rush-in of ions greatly increases the $Na^{+}$ concentration in the dynamic layer, naturally increasing the number of $Na^{+}$ ions passing through the channel. That increase alone creates the illusion of the increased $Na^{+}$ permeability. Of course, the ions moving across the segments change the field contributions, indirectly affecting the balanced state of the other ions. The permanently open ion channels only passively participate in the process, and their apparent penetrability changes with the changing fields. To provide channel penetrability data is meaningful only if ion concentrations and membrane thickness data are also available. The driving forces of the ions are different, and they produce, depending on the actual concentrations and actual local potential, the selectivity of the channels.
2.11.4 Ion pumps
"These pumps differ from ion channels in two important details. First, whereas open ion channels have a continuous water-filled pathway through which ions flow unimpeded from one side of the membrane to the other, each time a pump moves an ion, or a group of a few ions, across the membrane, it must undergo a series of conformational changes. As a result, the rate of ion flow through pumps is 100 to 100,000 times slower than through channels. Second, pumps that maintain ion gradients use energy, often in the form of ATP, to transport ions against their electrical and chemical gradients. Such ion movements are termed active transport." [1], page 101. We add that, finally, the energy source of the ion channels is also ATP. However, the potential energy acquired by the condenser (in the form of creating resting potential) is prepared "globally" (not at the proximity where it is used), and during the full life cycle of neuronal operation, mainly in the recovery period.
The role of ion pumps must also be revisited. It is not clear what a force can "transport ions against their electrical and chemical gradients", and how can they deliver the ions in sufficiently short times needed for the operation. In the resting state, instead of some magic lipid mechanism (ion pumps), the resultant potential (the sum of the electric and thermodynamic potential for the individual ions) moves the ions through the channels in the membrane. The resulting conductance on those channels is about 50 times smaller [10, 12] than that of the AIS. In a typical mammalian cell, see Table 1, the sum of the electric and thermodynamical potentials is $32\ [mV]\ K^{+}$ ions and $4\ [mV]\ Na^{+}$ ions. The same values for squids are $33\ [mV]\ K^{+}$ and $13\ [mV]\ Na^{+}$. Correspondingly, in the resting state, the resulting force moves sodium ions out of the cell and potassium ions into the cell (for more numeric examples see section 4.1). The so called "Na-K pump" is supposed to actively transport (using energy from ATP) sodium ions out of the cell and potassium ions into the cell. However, no mechanism is provided for how the ion pumps gain energy from ATP and forward it.
As Figure 6 in [15] displays, the ratio of $K+$ and $Na+$ changes sharply during the transient state. Theoretically, they assume that the neuron “pumps 3 $Na^{+}$ ions out of the cell and two potassium ions in”; experimentally, they show in their Fig. 6 that the ratio changes between 0.01 and 7.5. According to our theory, the magnitude of the current (and so the exchange ratio) depends on the resulting potentials, so the ratio must change as the membrane potential changes during AP. The classical theory cannot explain this behavior and leads to exciting conclusions: "Furthermore, we analyzed energy properties of each ion channel and found that, under the two circumstances, power synchronization of ion channels and energy utilization ratio have significant differences. This is particularly true of the energy utilization ratio, which can rise to above 100% during subthreshold activity." [15]
Our model says that both the concentration (due to $Na^{+}$ rush-in) and local membrane potential drastically change during the transient state. As long as the resulting potential is above $\approx 33\ mV$, the driving force acting on potassium ions is positive, i.e., the pump will move both ions out of the cell. The potential gradient near the membrane also explains the energy delivery: ATP, by hydrolysis, generates ions which move to the "plates" of the condenser. (as we discuss in section 5, in the transient state, the "setpoint" of the control circuit is also changed, and to restore the condenser’s voltage, those ions must be collected.)

### v3

@@ 4.7 The dynamic layers
This layer’s importance and operation are in line with the Earth’s atmosphere. Its features drastically deviate from the features of the bulks on its two sides. It is separated by a sharp contour on one side and an ill-defined border on the other; its volume is far from being homogeneous. Gravity keeps it in place, and it is at rest. However, sometimes, for some periods, also other (thermodynamic and electrical) forces are evoked inside it and lead to transient changes, moving huge masses with high speeds inside it. Their thickness and mass are negligible compared to those of the bulks on their two sides, and we can describe the bulks without considering their density, mass, size, etc. Still, this thin layer is responsible for the weather. It can temporarily absorb products of slow processes (water evaporation) and deliver masses of high density (much above its density, such as water, sand, etc.) to continental distances, creating the illusion that it temporarily stores that matter. Minor changes (natural ones, such as a slight difference in air temperature, and artificial ones, such as injecting condensation nuclei in clouds) can result in enormous changes. We can even imagine volcanic eruptions as semipermeable gates for material with apparently random operation and distribution of the injected material.
@@
The same is true of neuronal membranes and neuronal function. Now, we are at the point where their decades-old static description is insufficient. We need to derive the corresponding laws of motion to describe the neuron’s dynamic behavior. We need a meticulous and unusual analysis to derive them. In a neuron, in the abstraction science uses, we put together only an ionic solution, a semipermeable membrane, and currents that enter and leave it. All these belong to non-living matter. As experienced, at some combination of their parameters and gradients, qualitatively different phenomena happen, which, in the abstraction biology uses, are called signs of life; our system starts to belong to living matter. Given that the approximations, the derived abstractions, and the mathematical formalisms describing them are different for the two cases, it looks like we have two different, only loosely bound worlds. We realize we have arrived at the boundary of non-living and living matters, and we must go back to the first principles of science to clarify where their boundary is. However, by using our approach, we may defy that "the emergence of life cannot be predicted by the laws of physics" [46]. Our artificial duck looks, quacks, and swims like a duck.
@@ 4.8 Ionic currents in a cell
Notice that if any of the factors is zero, the macroscopic current $I$ is zero. Microscopic carriers must be present in the volume, and have charge, the cross section must not be zero, and the charge carriers must move with a potential-assisted speed, which needs an electrical field generated by the system, see Eq.(7). When there is no constraint force, the thermodynamic and electrical forces will produce an effective force that moves the ions. That is, the ions can move without an external electrical field, as experienced in biological systems. One of the fundamental mistakes by HH was neglecting the Coulomb force in ion-electrical interactions and assuming that diffusion processes or ideal electrical batteries create the electrical potential.
In biological systems, where ions represent the "current", it is either a native current (without an external potential) or an artificially injected current or potential generating a current. This way, the current is consistently accompanied by a change in concentration gradient, as the moving ions represent both mass and charge transfer simultaneously (also, the repulsion leads to a "skin effect": in electrical wires and axons, the charge travels near the outer surface). The potential and current are connected through the features of the medium (material) that hosts our measurement.
@@
(drag force) acting on it. A (microscopical) electrical force field $\frac{dV}{dz}$ inside the wire would accelerate the charge carriers continuously
@@
with a constant speed $v$. It is not the drift speed because of the electrical repulsion, it is a potential-assisted or potential-accelerated speed that can be by orders of magnitude higher. The medium in which the charge moves shows a (macroscopic, speed-dependent) counterforce $F_{d}$, which in steady state equals $F_{electric}$, that is :
@@
that is, for living matter
$E(z)$ $=\frac{\Delta V}{\Delta z}$ $in\ linear\ potential$ (24)
$=\frac{dV}{dz}$ $in\ graded\ potential$ (25)
$=E_{electric}+E_{(thermal)}(+E_{(external)})$ $in\ electrolytes$ (26)
The amount of current in a wire is not only influenced by the electric force field (specific resistance) but also by the number of charge carriers $n$. While the latter is commonly considered constant and part of the former, this is not necessarily the case for biological systems with electrically active structures inside. The medium’s internal structure ("the construction) introduces significant modifications. Applying an electric field to a biological wire can generate varying amounts of current as the number of charge carriers changes due to biological effects. For axons, we use a single-degree-of-freedom system, a viscous damping model, so the ions will move with a field-dependent constant velocity in the electrical field; so it takes time while they appear on the membrane. The activity of potential-controlled ion channels in its wall may change $n$ in various ways; furthermore, that change can result in ’delayed’ currents during measurement, for example, in clamping, as physiological measurements witness.
@@
The higher the resulting space derivative (gradient) and the fewer ions that can share the task of providing a current, the higher the speed. We hypothesized (it needs a detailed simulation) that in the case of this charged fluid, the electrical repulsion plays the role of ’viscosity’ [31]. The higher the charge density, the stronger the force equalizing the potential; so $\eta$ is the lower, the higher the charge density (proportional to $C_{k}$). For the sake of simplicity, we assume that the speed is proportional to the space gradient of the local concentration and potential. Recall that our equations refer to local gradients only. The electrical gradient can propagate only with the speed of the concentration gradient, given that only the chemically moved ions can mediate the electrical field [31]. The lower interaction speed limits the other interaction speed if the interactions generate each other.
@@
We introduce a time-dependent speed gradient $v(z,t)$ which plays a vital role in the processes that occur in living matter.
It is important to remember for alternating current experiments: the different ions will move at different speeds under the effect of the same $E(z,t)$, which is the function of the local concentrations and the frequency of the alternating current. The conduction speed sensitively depends on the temperature, and through it, the shape parameters [47] of the AP and especially the length of the "relative refractory" period [48].
@@ 4.9 Imitating slow currents
At the dawn of finding methods for describing neuronal operation, HH published high-precision measurements [14] enabling detailed testing of theories explaining the observed physiological behavior. Their good physical model that “movement of any charged particle in the membrane should contribute to the total current” lacked considering the mutual repulsion and finite speed of those particles (they are about 50,000 times heavier and a million times slower than the electrons, and they move in a closed space segment). Furthermore, they have started from the commonly used wrong assumption that conductance is a primary electrical entity. They could "find equations which describe the conductances with reasonable accuracy and are sufficiently simple for theoretical calculation of the AP and refractory period". The "reasonable accuracy" and the need for a simple calculation method (performed on a mechanical calculator) obscured the fact that the laws for electrons are not the same as the laws derived for ions.
@@
There is no voltage-dependent conductance [49]. Instead, the finite speed of ions and the wrong assumption that conductance is a primary entity mislead physiological research. With wording that "conductance changes", one states that charge carriers appear/disappear/reappear; that is, the charge and mass conservation are not fulfilled. Claims such as "the current cannot disappear, it has to go somewhere" [34], page 9, are wrong: there is no conservation law for current, which is the derivative of charge with respect to time. The correct physics background behind the early phenomenon [50] is that the number of charge carriers, at the cross section where the measurement is carried out, changes as a wave of ions with variable intensity moves slowly along the axon (they are apparently “created” for the static measurement arrangement). The changing conductance is a fallacy stemming from the idea of "fast current", which is valid for electrons but not for slow ions.
@@
During the regular operation of a neuronal membrane, after opening the ion channels, a vast amount of ions flows into the intracellular space from the extracellular space, somewhat similar to the effect of switching a clamping voltage. The essential difference is that the ions arrive through the axon to the joining point in clamping. In contrast, through the membrane’s ion channels, they directly contribute to the current on the membrane’s entire surface. The membrane’s size is finite, so with a finite current speed, it takes time until the charges on the membrane’s surface arrive at the AIS, as discussed for the axonal current. Fig. 5 shows the measured rising and falling edges of the clamping experiment performed by HH [14]. The continuous lines show the diagram lines fitted to the experimental data (notice again that the fitted polynomial significantly differs from the correct exponential function around the time 0.5 msec); the dashed line shows the expected result of the rising edge, calculated with the time constant of the falling edge. The number of charge carriers follows a saturating curve, which decreases the current in the axon. HH measured that their time constants differ far outside their experimental error, which would not be possible if they were the result of the charge-up and discharge of the same oscillator circuit. The measured rising edge corresponds to a charge-up process where the charge-up current is a saturating curve due to the change in the number of charge carriers in the axon; see Eq.(23).
Figure 5: The measured current, instead of conductance, (Fig. 2 from [14]), demonstrating that two different physical effects result in the rising and falling edge of the measured current.
@@ 4.10 Effects of membrane’s electricity
4.10.1 Ion channels
  • the potential generated across the membrane is entirely neglected
  • the ions have no driving force, and the hypothesized protein carriers are too slow [41]
  • the considered van der Waals force is too weak to be noticed by the ions
  • even if the weak van der Waals force would work for a single ion, the next ion would be rejected by the strong Coulomb force due to the first ion
  • the assumed force by the ’cation-attractive negative ends’ destabilizes the ion path: as the deviation from the central path increases, so does the deviating driving force
@@ 4.10.1 Ion channels
Our description is dynamic. The celebrated electron microscopic images are static: they capture only a snapshot of the ion channel and the ions (the ion channel is frozen). However, they cannot reveal how the transfer occurs. In Fig. 4, an ion channel is depicted in the middle of the figure with a diameter of about 3nm3\ nm for visibility. Furthermore, we assume that the ion’s size and, correspondingly, the thickness of the atomic layer in the electrolyte on the surface of the membrane is approximately 0.1nm0.1\ nm (although with their electrical double layers, they can also reach a size of 1nm1\ nm). For comparison, recall that an Atomic Force Microscope (AFM) tip’s size is typically about 10nm10\ nm, and the size of the tip of the clamp pipette is in the range of 1,000nm1,000\ nm and the size of the soma in the range of 10,000nm10,000\ nm.
@@
Figure 4 also hints at how ions can move in the proximity of the membrane. On the left side, the electrical field increases toward the membrane, so ions cannot move in that direction without an initial force. The case is the same on the right side because of the opposite charge. An ion must gain energy to get closer to the membrane, which means, at the same time, higher potential energy. At that position, it is attracted by the charges on the opposite side and kept back by the mechanical counterforce of the membrane plate on this side. This situation explains why the ions do not diffuse into the bulk region. (The charges compress the membrane that changes the membrane’s capacity.) The ion channels can work in continuous mode (the "resting ion channels") and impulse mode (the "transient ion channels" are gated by "caps"): "Two types of ion channels—resting and gated— have distinctive roles in neuronal signaling. Resting channels are primarily important in maintaining the resting membrane potential, the electrical potential across the membrane in the absence of signaling. Some types of resting channels are constitutively open and are not gated by changes in membrane voltage; other types are gated by voltage but can open at the negative resting potential of neurons. Most voltage-gated channels, in contrast, are closed when the membrane is at rest and require membrane depolarization to open.[17], page 126. Closing ion channels is important for making ungraded layer potentials.
@@
The resulting electrical field cannot accelerate an ion except when the ion is at the beginning of the open ion channel across the membrane. Here, the mechanical counterforce is missing, and the field drops to the value generated by the condenser inside the plates. Here, a vast electrical field accelerates the ions to high speed (although the force required to overcome the viscosity reduces the electrostatic acceleration). The empty red circles show an ion that traverses through the channel. The figure shows the electrical fields on both sides of the channel at the beginning and end. (On the right side, the electrical field is negative, but so is the charge of the ions.)
@@
The electrical acceleration sharply decreases after leaving the ion channel (see the right side). As shown by the difference in the vertical positions of the filled red and green balls, the difference in the electrical field values is huge across the membrane ($E_{accelerate}$) and considerable inside the dielectric electrolyte ($E_{assist}$), as shown by the difference in the electrical field values of the green balls. The field still accelerates, but significantly less intensely than the one in the gap. Suppose we assume that the ion quickly accelerates to its Stokes-Einstein speed. In that case, the travel speeds in those space regions are proportional to the difference of the fields in the shown positions and the reciprocals of their travel times.
@@
The passing ions decrease the concentration and the electrical field on the high-concentration side, and, conversely, increase them on the low-concentration side. At the exit, the field continues to accelerate the ion. However, the increasing potential and concentration increasingly decelerate the passing ions, so they quickly brake them to the assisted speed. Then, the effect of the electrical field cancels, and the ions will move with their diffusion speed ("the ion stops"). The local potentials, connected to the operation of the ion channels, act as an "electrical lock".
@@
Moving a charged particle inside the ion channels (nanopores) in membranes is a complex phenomenon and its details are discussed in [27]. They have studied theoretically and experimentally the generation of membrane potential at zero current in nanoporous membranes separating two reservoirs with different electrolyte concentrations. Although not specifically for our environment, they confirmed our statement that the membrane potential is of electrostatic origin (exists at zero current) and its value is in the order of a couple of $mV$. Their finding that "the corresponding Donnan potentials appear at the pore entrance and exit, leading to a dramatic enhancement of membrane potential in comparison with an uncharged dielectric membrane" inspires beginning further research in uncovering quantitatively the details of the operation of the dynamic layer governing neuronal operation; furthermore, moving toward studying non-equilibrium processes.
@@
We want to describe the change of voltage due to limited resources using an idea similar to the predator-prey model. We have charged sheets (represented by ions in the electrolyte) with potential $U_{H}$ and $U_{L}$ on the two surfaces of the membrane, plus we consider the corresponding neighboring layers on their side toward the "bulk" of the segment. We assume the layers’ capacities are constant and equal, so the change in charge is linearly proportional to the change in voltage.
@@
We consider that a charge transfer happens from the layer with potential $U_{H}$ to the layer with potential $U_{L}$  (that is, the same charge is removed and added, respectively) with a high speed (we call it potential-accelerated speed); furthermore, that with a much lower speed (we call it potential-assisted speed) the charge in the neighboring layer increases and decreases, respectively, and those voltages are ($U_{H-1}$ and $U_{L+1}$).
@@
We assume constant layer capacity and that the amount of transferred charge (or voltage) is proportional to the voltage difference between the layers.
@@
As Fig. 7 depicts, voltages $U_{H}$ and $U_{L}$ quickly approach their balanced values. When they get equal, the driving force cancels. In the meantime, the voltages $U_{H-1}$ and $U_{L+1}$ tend to approach $U_{H}$ and $U_{L}$, respectively. The diffusional and energy-producing processes, depending on the local conditions, change the voltages on the two ends of the channel, creating the illusion that the channel opens and closes (ad-hoc constants are used to illustrate the concept). As [35] analyzed, this on/off behavior of currents can be observed for natural and synthetic lipid membranes, without caps. Despite the slightly different experimental conditions, the amplitudes and typical time scales are similar.
4.10.2 Maxwell demon
After all, gated ion channels, see Fig. (4), operate as Maxwell-demon’-like objects built into the separating membrane (from the point of view of the segments and the observer). In the classical theory, some power opens them, and they autonomously transfer ions in a potential-accelerated operating mode, and then that power puts the cap back on the channel. The segments are separated by a semipermeable membrane with ion channels capped on its surface. As long as the caps of the ion channels are closed and the ion concentrations on both sides of the membrane are the same, we do not observe any change in the state of the solutions. Although the channels can stochastically open, close, and re-open, they transmit charge quanta of varying degrees of definition. Even the channels can recognize the ions’ chemical nature and transmit only a selected ion type. The channels are passive during those processes, although the enormous voltage gradient can rearrange their structure and change their behavior through that. The demons also coordinate their actions using the layer containing charges as a communication medium; their population maintains a well-defined macroscopic current across the membrane. Physics can also explain the magic behind their operation.